OpenStudy (anonymous):

suppose f(x) = x+1 and g(x)= x^2-4x . find the inverse function of (f o g)(x), x >/= 2.

OpenStudy (anonymous):

so the equation would be (fog) (x)= (x^2-4x)+1

OpenStudy (anonymous):

h(x)=(f o g)(x)=x^2-4x +1 x^2-4x +1=h (x-2)^2+1 -4 =h (x-2)^2=h+3 x=+-sqrt(h+3) +2

OpenStudy (anonymous):

I have the answers, but not the solutions.. the answer to this one is 2 + sqrt x+3 , so I think that 2 goes in front of the sqrt?

OpenStudy (anonymous):

but otherwise, thanks. I couldn't figure out what to do with 2 exponents

OpenStudy (anonymous):

ohh, right dindt see " x >/= 2". So only positive sqrt is needed.

OpenStudy (anonymous):

you wellcome